MEASUREMENT OF THE PURE ROTATIONAL QUASIBOUND SPECTRA OF HeH^{+} AND HeD^{+} IN A LABORATORY PLASMA BY DIRECT LASER ABSORPTION

\underline{Z. LIU}, W. Y. FAN and P. B. DAVIES, Department of Chemistry, Uiversity of Cambridge, Cambridge CB2 1EW, UK.

The pure rotational quasibound spectra of HeH^{+} and HeD^{+} have been observed in the mid infrared region using diode laser velocity modulation absorption spectroscopy. Four quasibound to quasibound(Q−Q) and three bound to quasibound(B − Q) transitions of HeH^{+} and one bound to quasibound transition of HeD^{+} have been detected in a low pressure ac glow discharge. The highest quasibound level involved in the spectra (HeH^{+}, v=0, J=26) is 2000 cm^{−1} above the dissociation limit. The measured transition frequencies are in very good agreement with ab intio calculations. Two of the Q − Q transitions have wider linewidths due to lifetime broadening.